# Understanding the Potential of FPGA-Based Spatial Acceleration for Large Language Model Inference

### 23 Dec 2023 v1

@@ 3.2. Resource Constraints
The core computational element for linear operators is the MAC unit. Let $M_{i}$ denote the compute power, in terms of the number of MACs per cycle allocated to each matrix multiplication kernel, where $i$ ranges over $q$, $k$, $v$, $a_{1}$, $a_{2}$, $p$, $f_{1}$, and $f_{2}$, based on the notation in Table 1. We quantize the matrix multiplication to integer inputs for maximum efficiency, which has been proven to be effective by many recent studies  (Xiao et al., 2023; Dettmers et al., 2022; Shen et al., 2020; Kim et al., 2021). Quantization enables single-cycle accumulation. As a result, one multiply-accumulator (MAC) unit can provide a 1 MAC/cycle throughput with a properly pipelined multiplier. Therefore, the latency for the $Q$ projection can be calculated as $ld^{2}/M_{q}$ cycles, considering that the total number of MACs computed in this operator is $ld^{2}$.
@@
where $M_{\text{tot}}$ represents the total available compute power of an FPGA in terms of MACs per cycle, which can be obtained from the official data sheets. For FPGAs with specialized compute blocks (Xilinx, 2022c; Langhammer et al., 2021), we can convert their compute power to match the frequency of the programming logic, thus obtaining an overall value $M_{\text{tot}}$ for the entire FPGA.
@@
Parameter buffers. To optimize an FPGA-based dataflow design, we assume that all the quantized parameters can be accommodated in on-chip or off-chip memory. Suppose all the linear weights are quantized to $b_{W}$ bits, and the size of the linear operator $i$ is $s_{i}$. The total size of the buffers is $S_{param}=\sum_{i\in\{q,k,v,p,f_{1},f_{2}\}}s_{i}b_{W}=(4d^{2}+2dd_{\text{FFN}})b_{W}$ if storing on-chip. If the parameters are too large to fit in on-chip memory, we can store the parameters in DRAM and tile the parameters with size $M_{i}$ on-chip, then the total tiled buffer size is $S_{tile}=\sum_{i\in\{q,k,v,p,f_{1},f_{2}\}}M_{i}b_{W}$. To hide the memory access latency, we need to double buffer those parameters, so the final buffer size of the $i$-th linear operator is $2S_{tile}$.
KV Cache. When conducting matrix multiplication, at least one of the matrices’ elements must be accessed repeatedly so that a buffer is required. Given that parameters are already buffered, only the SDP requires buffering for at least one of the input matrices. In our case, we choose to buffer $K$ and $V$, which will be later passed to the decode stage as the KV cache. We also double buffer $K$ and $V$ matrices to improve throughput. The final buffer size is $S_{KV}=4l_{\max}db_{A}$, where $b_{A}$ is the bitwidth of the activation and $l_{\max}$ is the maximum sequence length supported by the model.
FIFOs. The intermediate results between linear operators flow in FIFOs since the linear operators sequentially access them. For the initial residual connection, we assume that the input tensors are fetched from off-chip memory to obviate the need for additional buffering. However, for the second residual connection related to the FFN, it is necessary to use an intermediate buffer to store the projection’s activation $X_{act}$ before the FFN. This buffer simultaneously serves as a bypass path. To avoid deadlock, the buffer must possess sufficient capacity to store $X_{act}$. We simply create a FIFO of size $ldb_{A}$ to store it. For other FIFO connections, we assume a FIFO depth of $s$ and one FIFO connecting each layer in Figure 2, so the total FIFO size is equal to $S_{FIFO}=16sb_{A}+ldb_{A}$.
@@
(3) $S_{param}C$ $<DRAM_{\text{tot}}\,,$
$\sum S_{i}C$ $<SRAM_{\text{tot}}\,,i\in\{tile,KV,FIFO\}\,,$
@@
If the parameters are stored off-chip, we need to consider the impact of off-chip memory bandwidth. We can assume there are enough on-chip buffers to hide the off-chip memory access latency and convert a DRAM-friendly data layout to a PE-friendly one. Similar to §4.2.2, we use $r_{i}$ to denote the data reuse factor of a linear operator with $M_{i}$ MAC units. Effectively, $M_{i}/r_{i}$ weights must be loaded from off-chip memory per cycle to feed the MAC units, requiring a bandwidth of:
@@
where $freq$ is the achieved frequency of FPGA. If the total required bandwidth, $\sum_{i}C{B_{i}}(i\in\{q,k,v,p,f_{1},f_{2}\})$, exceeds the maximum device bandwidth, the inference becomes bandwidth bound.
@@ 3.3. Performance Estimation
We construct the pipeline diagram as shown in Figure 4. As mentioned in §3.2.2, since we need to store the $K$ and $V$ values after the linear operators, there is an implicit synchronization point between the $q$/$k$/$v$ operator and the latter SDP and FFN parts. The computation of them cannot be overlapped. Notice the $q$/$k$/$v$ operator can be performed in parallel since they do not have any dependencies. After the $k$ and $v$ have been fully calculated, the latter computation of SDP and FFN can be greatly overlapped, since they do not need to wait for all the results to perform the next operator, and the results of the previous operator can be directly streamed into the next operator as input. Moreover, since different Transformer layers share the same architecture, their computation can also be overlapped without waiting for the result of the previous layer.
@@
(10) $\frac{ld^{2}}{M_{q,k,v,p}}=\frac{ld^{2}/h}{M_{a_{1},a_{2}}}h=\frac{ldd_{\text{FFN}}}{M_{f_{1},f_{2}}}$
@@ 3.4. Distributed Inference
Various methods exist for facilitating inter-FPGA communication, including communication through the host, PCI-E Peer-to-Peer (P2P), and on-device Ethernet. We mainly consider the third approach since it does not necessitate orchestration from the host and provides higher bandwidth compared to other alternatives. For example, the AMD Alveo U280 FPGA provides two QSFP ports (Xilinx, 2022b), each capable of carrying 100 Gb/s Ethernet data over optical fibers, which ensures robust and high-speed inter-FPGA communication. Most of the time, we cannot fully utilize the network bandwidth and need to pay for the package header overheads. Suppose the theoretical network bandwidth between two FPGA devices is $B$ bits per second (bps), and the efficiency of the network is $\alpha$, so we can have the effective bandwidth as $\alpha B$, where $\alpha$ can be obtained through network benchmarking.
@@
As mentioned in §2.2, we have various parallelization schemes when considering multiple devices. We first analyze tensor parallelism (TP). Traditionally, TP requires two all-reduce operations within one Transformer layer as depicted in Figure 3. However, these communicative operations are implemented in a blocking way. Figure 5(a) shows the subsequent FFN module needs to wait for the completion of the all-reduce process before it can conduct computation (Wang et al., 2023). Notice that the all-reduce operation only involves fetching results from other devices and adding the result to its local tensor. Given that the output of MHA is a sequential stream, we can perform elementwise addition in a non-blocking manner. As soon as the kernel receives enough data, it can initiate data transfer to other devices without waiting for the remaining data to be computed. This leads to substantial synchronization time savings as shown in Figure 5(b).
@@
As we have already implemented dataflow inside a device, pipeline parallelism (PP) essentially extends the dataflow to $p_{2}$ devices with a tensor of size $ld$ communicated in between. Here we only split the pipeline between two Transformer layers. Therefore, the latency of Equation (8) becomes
@@
where $p_{1}$ and $p_{2}$ are the size of a TP group and a PP group (Shoeybi et al., 2019). Additionally, the memory requirements of Equations (3) and (5) need to be divided by $p_{1}$ to satisfy the constraints of multiple devices.
Notice we only discuss two basic parallelism schemes for Transformer models. Some recent works may partition the sequence dimension and leverage reduce-scatter and all-gather to reduce the overheads of all-reduce (Narayanan et al., 2021b; Korthikanti et al., 2022). The communication time can be similarly analyzed, and we will not discuss them here. The optimal parallelism scheme on multiple devices (Pope et al., 2023; Unger et al., 2022; Zheng et al., 2022; Miao et al., 2022; Xie et al., 2022) is out of the scope of this paper, and we will leave it as future works.
@@ 4.1. Overview of Workloads and Hardware
Table 2 lists several widely used models that we choose for performance estimation. Such models include BERT-base (Devlin et al., 2018), a representative encoder-based model, GPT2 (Radford et al., 2019), the only open-sourced model in the GPT family, LLaMA-7B (Touvron et al., 2023a, b), an open-sourced model trained by Meta, and Vicuna-13B (Chiang et al., 2023), the best non-commercial model on Chatbot Arena (Zheng et al., 2023).
@@
  BERT (Devlin et al., 2018) Encoder 110M 12 12 768 3072
GPT2 (Radford et al., 2019) Decoder 355M 24 16 1024 4096
LLaMA2 (Touvron et al., 2023b) Decoder 7B 32 32 4096 11008
Vicuna (Chiang et al., 2023) Decoder 13B 40 40 5120 13824
To see the performance differences between FPGAs and GPUs, we pick and list several representative devices in Table 3. For FPGAs, Alveo U280 and Agilex 7 are two FPGAs that are widely used in cloud servers but not specially optimized for AI workloads; Versal VCK5000 and Stratix 10 NX FPGAs are designed for accelerating AI applications with specialized hardware units such as AI Engine (AIE) (Xilinx, 2022a) or Tensor Blocks (Langhammer et al., 2021). Versal VHK158 is the latest FPGA with HBM2e released by AMD in 2023. For GPUs, RTX 2080Ti is a high-end GPU designed for personal usage with a similar process node and release date to U280; A100 is the most deployed GPU in data centers to conduct LLM training and inference (OpenAI, 2023).
@@
Alveo U280 (Xilinx, 2021) Versal VCK5000 (Xilinx, 2022c) Versal VHK158 (Xilinx, 2023) Stratix 10 NX 2100 (Langhammer et al., 2021) Agilex 7 AGM039 (Intel, 2022) GeForce RTX 2080 Ti Tesla A100
@@ 4.2. Single-Device Performance Estimation
We first conduct experiments on a single device with pre-trained BERT-base and GPT2 models from HuggingFace Hub (Wolf et al., 2019). For GPU baselines, we run the models with PyTorch 2.0 (Paszke et al., 2019) and CUDA 11.7, and measure the performance on both RTX 2080Ti and A100 GPUs. The host machine runs an Intel Xeon Silver 4114 CPU at 2.20GHz with 40 cores. We follow the common practice to measure the out-of-the-box FP16 performance (Aminabadi et al., 2022; Nvidia, 2022; Radford et al., 2019).
@@
From Figure 6(a) and (b), we can see 2080Ti and A100 GPUs maintain an almost constant curve, which means changing sequence length does not significantly impact their performance. However, since FPGAs are blocked by the limited compute and memory resources on-board, the latency increases linearly in terms of sequence length as described in Equation (8), and demonstrates significantly longer latency when the sequence length is large (e.g., 512). It is worth noting that even when making use of state-of-the-art AI-optimized FPGAs like the Stratix 10 NX and Versal VCK5000, the situation does not improve significantly. This is because these modern FPGAs are equipped with DDR or older versions of HBM, which makes parameter loading from off-chip memory the bottleneck. Even though these FPGAs have highly efficient computation blocks, the compute units have to wait for the memory to fetch data, resulting in suboptimal performance.
On the contrary, when the sequence length is small, GPUs suffer from underutilizing the compute resources, and FPGAs can achieve a significantly lower latency compared to GPUs. This is particularly evident in the decode stage, as shown in Figure 6(c). The latency of RTX2080 and A100 GPUs increases since the memory access takes control compared to computation in the decode stage. Although FPGAs are also bounded by off-chip memory bandwidth, they can always perform better than GPUs since the computation is small when the sequence length is equal to one. Therefore, FPGAs can easily achieve GPU-level performance even with a small $M$.
@@
Lastly, we investigate how quantization impacts the required memory bandwidth. As shown in Figure 8(c), the low-bit weight quantization can significantly alleviate the demands of off-chip memory access. By reducing the volume of data needed in each cycle, it can achieve a larger compute power $M$, thus leading to a better performance. In particular, quantizing the model to a 2-bit representation yields a performance boost exceeding an order of magnitude when compared to a 16-bit weight quantization scheme. Recent research (Zhao et al., 2023; Chee et al., 2023) has demonstrated that 4-bit or even 2-bit quantization can be implemented without compromising model accuracy, which makes efficient LLM deployment on FPGAs possible.

### 07 Apr 2024 v2

@@ 3.2. Resource Constraints
The core computational element for linear operators is the MAC unit. Let $M_{i}$ denote the compute power, in terms of the number of MACs per cycle allocated to each matrix multiplication kernel, where $i$ ranges over $q$, $k$, $v$, $a_{1}$, $a_{2}$, $p$, $f_{1}$, and $f_{2}$, based on the notation in Table 1. We quantize the matrix multiplication to integer inputs for maximum efficiency, which has been proven to be effective by many recent studies  (xiao2023smoothquant, ; dettmers2022llmint8, ; sheng2020qbert, ; kim2021ibert, ). Quantization enables single-cycle accumulation. As a result, one multiply-accumulator (MAC) unit can provide a 1 MAC/cycle throughput with a properly pipelined multiplier. Therefore, the latency for the $Q$ projection can be calculated as $ld^{2}/M_{q}$ cycles, considering that the total number of MACs computed in this operator is $ld^{2}$.
@@
where $M_{\text{tot}}$ represents the total available compute power of an FPGA in terms of MACs per cycle, which can be obtained from the official data sheets. For FPGAs with specialized compute blocks (e.g., AI Engine (vck5000, ) and AI Tensor Blocks (stratix10, )), we can convert their compute power to match the frequency of the programming logic, thus obtaining an overall value $M_{\text{tot}}$ for the entire FPGA. For example, the VCK5000 FPGA (vck5000, ) has 400 AI Engines, each of which can compute 128 MACs/cycle at 1GHz. Therefore, the equivalent compute power at 250MHz is 128$\times$400$\times$1GHz/250MHz, which is 204800 MACs/cycle.
@@
Parameter buffers. To optimize an FPGA-based dataflow design, we assume that all the quantized parameters can be accommodated in on-chip or off-chip memory. Suppose all the linear weights are quantized to $b_{W}$ bits, and the size of the linear operator $i$ is $s_{i}$. The total size of the buffers is $S_{\text{param}}=\sum_{i\in\{q,k,v,p,f_{1},f_{2}\}}s_{i}b_{W}=(4d^{2}+2dd_{\text{FFN}})b_{W}$ if storing on-chip. If the parameters are too large to fit in on-chip memory, we can store the parameters in DRAM and tile the parameters with size $M_{i}$ on-chip, then the total tiled buffer size is $S_{\text{tile}}=\sum_{i\in\{q,k,v,p,f_{1},f_{2}\}}M_{i}b_{W}$. To hide the memory access latency, we need to double buffer those parameters, so the final buffer size of the $i$-th linear operator is $2S_{\text{tile}}$.
KV Cache. When conducting matrix multiplication, at least one of the matrices’ elements must be accessed repeatedly so that a buffer is required. Given that parameters are already buffered, only the SDP requires buffering for at least one of the input matrices. In our case, we choose to buffer $K$ and $V$, which will be later passed to the decode stage as the KV cache. We also double buffer $K$ and $V$ matrices to improve throughput. The final buffer size is $S_{\text{KV}}=4l_{\max}db_{A}$, where $b_{A}$ is the bitwidth of the activation and $l_{\max}$ is the maximum sequence length supported by the model. Notice KV cache can also be tiled on-chip, which can leverage a similar analysis above.
FIFOs. The intermediate results between linear operators flow in FIFOs since the linear operators sequentially access them. For the initial residual connection, we assume that the input tensors are fetched from off-chip memory to obviate the need for additional buffering. However, for the second residual connection related to the FFN, it is necessary to use an intermediate buffer to store the projection’s activation $X_{\text{act}}$ before the FFN. This buffer simultaneously serves as a bypass path. To avoid deadlock, the buffer must possess sufficient capacity to store $X_{\text{act}}$. We simply create a FIFO of size $ldb_{A}$ to store it. For other FIFO connections, we assume a FIFO depth of $s$ and one FIFO connecting each layer in Figure 2, so the total FIFO size is equal to $S_{\text{FIFO}}=16sb_{A}+ldb_{A}$.
@@
(3) $S_{\text{param}}C$ $<DRAM_{\text{tot}}\,,$
$\sum S_{i}C$ $<SRAM_{\text{tot}}\,,i\in\{\text{tile},\text{KV},\text{FIFO}\}\,,$
@@
If the parameters are stored off-chip, we need to consider the impact of off-chip memory bandwidth. Similar to §4.2.2, we use $r_{i}$ to denote the data reuse factor of a linear operator with $M_{i}$ MAC units. Effectively, $M_{i}/r_{i}$ weights must be loaded from off-chip memory per cycle to feed the MAC units, requiring a bandwidth of:
@@
where $freq$ is the achieved frequency of FPGA. If the total required bandwidth, $\sum_{i}C{B_{i}}(i\in\{q,k,v,p,f_{1},f_{2}\})$, exceeds the maximum device bandwidth, the inference becomes bandwidth bound. Notice this bandwidth requirement needs to be analyzed for each operator individually if the data loading requires accessing multiple DDR or HBM channels.
@@ 3.3. Performance Estimation
We construct the pipeline diagram as shown in Figure 4. As mentioned in §3.2.2, since we need to store the $K$ and $V$ values after the linear operators, there is an implicit synchronization point between the $q$/$k$/$v$ operator and the latter SDP and FFN parts. The computation of them cannot be overlapped. Notice the $q$/$k$/$v$ operator can be performed in parallel since they do not have any dependencies. After $k$ and $v$ have been fully calculated, the subsequent computations of SDP and FFN can be greatly overlapped. This is because these operations do not need to wait for all the results to perform the next operation. The results of the previous operation can be directly streamed into the next operation as input. Moreover, since different Transformer layers share the same architecture, their computation can also be overlapped without waiting for the result of the previous layer.
@@
(10) $\frac{ld^{2}}{M_{q,k,v,p}}=\frac{l^{2}d/h}{M_{a_{1},a_{2}}}h=\frac{ldd_{\text{FFN}}}{M_{f_{1},f_{2}}}$
@@ 3.4. Distributed Inference
Various methods exist for facilitating inter-FPGA communication, including communication through the host, PCI-E Peer-to-Peer (P2P), and on-device Ethernet. We mainly consider the third approach since it does not necessitate orchestration from the host and provides higher bandwidth compared to other alternatives. For example, the AMD Alveo U280 FPGA provides two QSFP ports (qsfp, ), each capable of carrying 100 Gb/s Ethernet data over optical fibers, which ensures robust and high-speed inter-FPGA communication. Most of the time, we cannot fully utilize the network bandwidth and need to pay for the package header overheads. Suppose the theoretical network bandwidth between two FPGA devices is $B$ bits per second (bps), and the efficiency of the network is $\alpha$, so we can have the effective bandwidth as $\alpha B$, where $\alpha$ can be obtained through network benchmarking.
@@
As mentioned in §2.2, we have various parallelization schemes when considering multiple devices. We first analyze tensor parallelism (TP). As shown in Figure 3, the parameters of the linear operations are partitioned across different devices. For example, suppose the weight parameters of the two FFN layers $f_{1}$ and $f_{2}$ are $A$ and $B$, then we can partition $A$ along its column and partition $B$ along its row, and obtain
$\sigma(ZA)B=\sigma\left(Z\begin{bmatrix}A_{1}&A_{2}\end{bmatrix}\right)\begin{bmatrix}B_{1}\\ B_{2}\end{bmatrix}=\sigma(ZA_{1})B_{1}+\sigma(ZA_{2})B_{2}\,,$
where $\sigma$ is the GeLU function. Therefore, apart from partitioning $A$ and $B$, we need to insert an all-reduce operation to aggregate the partial results on each device to ensure correctness. The partitioned parameters will be stored on different devices. For example, $A_{1}$ will be on the first FPGA, and $A_{2}$ will be on the second FPGA. A similar partition scheme can be applied for MHA, and we refer the readers to (shoeybi2019megatron, ) for more details.
Based on this partition scheme, TP requires two all-reduce operations within one Transformer layer. However, these communicative operations are implemented in a blocking way. Figure 5(a) shows the subsequent FFN module needs to wait for the completion of the all-reduce process before it can conduct computation (wang2023overlap, ). Notice that the all-reduce operation only involves fetching results from other devices and adding the result to its local tensor. Given that the output of MHA is a sequential stream, we can perform elementwise addition in a non-blocking manner. As soon as the kernel receives enough data, it can initiate data transfer to other devices without waiting for the remaining data to be computed. This leads to substantial synchronization time savings as shown in Figure 5(b).
@@
As we have already implemented dataflow inside a device, pipeline parallelism (PP) essentially extends the dataflow to $p_{2}$ devices with a tensor of size $ld$ communicated in between. Here, we only split the pipeline between two Transformer layers so the results of the previous device can be directly streamed to the next device in the same PP group. Notice TP and PP can be combined to conduct model inference (narayanan2021megatronv2, ), and the latency of Equation (8) becomes
@@
where $p_{1}$ and $p_{2}$ are the size of a TP group and a PP group (shoeybi2019megatron, ). Additionally, the memory requirements of Equations (3) and (5) need to be divided by $p_{1}$ to satisfy the constraints of multiple devices.
Notice we only discuss two basic parallelism schemes for Transformer models. Some recent works may partition the sequence dimension and leverage reduce-scatter and all-gather to reduce the overheads of all-reduce (narayanan2021megatronv2, ; korthikanti2022reducing, ). The communication time can be similarly analyzed, and we will not discuss them here. The optimal parallelism scheme on multiple devices (pope2022googleinf, ; colin2022unity, ; zheng2022alpa, ; miao2022galvatron, ; xie2022optimalplacement, ) is out of the scope of this paper, and we will leave it as future works.
@@ 4.1. Overview of Workloads and Hardware
Table 2 lists several widely used models that we choose for performance estimation. Such models include BERT-base (devlin2018bert, ), a representative encoder-only model, GPT2 (radford2019gpt2, ), the only open-sourced model in the GPT family, LLaMA-7B (touvron2023llama, ; touvron2023llama2, ), an open-sourced model trained by Meta, and Vicuna-13B (chiang2023vicuna, ), the best non-commercial model on Chatbot Arena (zheng2023chatbotarena, ).
@@
  BERT (devlin2018bert, ) Encoder 110M 12 12 768 3072
GPT2 (radford2019gpt2, ) Decoder 355M 24 16 1024 4096
LLaMA2 (touvron2023llama2, ) Decoder 7B 32 32 4096 11008
Vicuna (chiang2023vicuna, ) Decoder 13B 40 40 5120 13824
To see the performance differences between FPGAs and GPUs, we pick and list several representative devices in Table 3. For FPGAs, Alveo U280 and Agilex 7 are two FPGAs that are widely used in cloud servers but not specially optimized for AI workloads; Versal VCK5000 and Stratix 10 NX FPGAs are designed for accelerating AI applications with specialized hardware units such as AI Engine (AIE) (xilinxAIE, ) or Tensor Blocks (stratix10, ). Versal VHK158 is the latest FPGA with HBM2e released by AMD in 2023. For GPUs, RTX 2080Ti is a high-end GPU designed for personal usage with a similar process node and release date to U280; A100 is the most deployed GPU in data centers to conduct LLM training and inference (openai2023gpt4, ).
@@
Alveo U280 (u280, ) Versal VCK5000 (vck5000, ) Versal VHK158 (vhk158, ) Stratix 10 NX 2100 (stratix10, ) Agilex 7 AGM039 (agilex7, ) GeForce RTX 2080 Ti Tesla A100
@@ 4.2. Single-Device Performance Estimation
We first conduct experiments on a single device with pre-trained BERT-base and GPT2 models from HuggingFace Hub (wolf2019huggingface, ). For GPU baselines, we run the models with PyTorch 2.0 (paszke2019pytorch, ) and CUDA 11.7, and measure the performance on both RTX 2080Ti and A100 GPUs. The host machine runs an Intel Xeon Silver 4114 CPU at 2.20GHz with 40 cores. We follow the common practice to measure the out-of-the-box FP16 performance (aminabadi2022dsinf, ; fastertransformer2022, ; radford2019gpt2, ).
@@
From Figure 6(a) and (b), we observe that the 2080Ti and A100 GPUs maintain an almost constant curve when the sequence length is less than 1024. This behavior is primarily attributed to the high kernel launch overheads for hundreds of CUDA kernels in PyTorch, which overshadow the computation time. However, when the sequence length exceeds 1024, GPUs become computation-bound, resulting in higher latency. In contrast, the latency of FPGAs increases linearly with the sequence length, as described in Equation (8), and demonstrates significantly longer latency when the sequence length is large (e.g., 512). It is worth noting that even when making use of state-of-the-art AI-optimized FPGAs like the Stratix 10 NX and Versal VCK5000, the situation does not improve significantly. This is because these modern FPGAs are equipped with DDR or older versions of HBM, which makes parameter loading from off-chip memory the bottleneck. Even though these FPGAs have highly efficient computation blocks, the compute units have to wait for the memory to fetch data, resulting in suboptimal performance. Moreover, many FPGAs struggle to achieve both high memory bandwidth and compute power simultaneously. Consequently, for VHK158, performance deteriorates when the sequence length reaches 4096, as it shifts from being memory-bandwidth bound to compute-bound. Further scaling the sequence length larger than 4096 may lead to out-of-memory for both FPGAs and GPUs, and only A100 can handle such large sequence lengths, so the latency is not plotted on the figure.
The decode stage is on the contrary, as shown in Figure 6(c). When the sequence length is small, GPUs suffer from underutilizing the compute resources, and FPGAs can achieve a significantly lower latency compared to GPUs. The latency of RTX2080 and A100 GPUs increases since the memory access takes control compared to computation in the decode stage. Although FPGAs are also bounded by off-chip memory bandwidth, they can always perform better than GPUs since the computation is small when the sequence length is equal to one. Therefore, FPGAs can easily achieve GPU-level performance even with a small $M$.
@@
Lastly, we investigate how quantization impacts the required memory bandwidth. As shown in Figure 8(c), the low-bit weight quantization can significantly alleviate the demands of off-chip memory access. By reducing the volume of data needed in each cycle, it can achieve a larger compute power $M$, thus leading to a better performance. In particular, quantizing the model to a 2-bit representation yields a performance boost exceeding an order of magnitude when compared to a 16-bit weight quantization scheme. Recent research (zhao2023atom, ; chee2023quip, ) has demonstrated that 4-bit or even 2-bit quantization can be implemented without compromising model accuracy, which makes efficient LLM deployment on FPGAs possible.